OpenStudy (anonymous):

Which of the following is the best estimate of 5 and seven-eighths minus 1 and one-tenth? 4 and one-half 4 3 and one-half 5

OpenStudy (wolf1728):

5 and 7/8 = 5.875 1 and 1/10 = 1.1 5.875 - 1.1 = 4.775

OpenStudy (wolf1728):

Seems 5 is the best answer.